OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation ((x-5)^2/81)-((y-1)^2)/144)=1??? help please :c

OpenStudy (anonymous):

\[\frac{(x-5)^2}{9^2}-\frac{(y-1)^2}{12^2}=1\] is a start the center is therefore \((5,1)\)

OpenStudy (anonymous):

you need to know the center in order to find the vertices you also need to know what it looks like do you know what this one looks like?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

because the \(x^2\) term comes first, it looks like this 

OpenStudy (anonymous):

you need to know it looks like this to see that the vertices are to the left and right of the center, not up and down

OpenStudy (anonymous):

general form is \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]and you have \[\frac{(x-5)^2}{9^2}-\frac{(y-1)^2}{12^2}=1\] so \(a=9\)

OpenStudy (anonymous):

this tells you the vertices are 9 units to the left and right of the center, which is \((5,1)\)

OpenStudy (anonymous):

im stilll confused :/ im sorry

OpenStudy (anonymous):

do you see tjat \(a^2=81\) so \(a=9\) ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok and do you see from the equation that the center is \((5,1)\) ?

OpenStudy (anonymous):

yes, but for the vertices would i just add and subtract 9 from 5?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is why you need the center first

OpenStudy (anonymous):

what else do you need? the foci?

OpenStudy (anonymous):

yeah the foci, and just to double check would the vertices be (14,1) and (-4,1)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for the foci, again they will be to the left and right of the center, but you need another number

OpenStudy (anonymous):

\[c^2=a^2+b^2\] just like pythagoras then in your case \[c^2=81+144=225\] and so \(c=15\) now go 15 units left and right for the foci

OpenStudy (anonymous):

okay, thanks so much!(:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

btw if you ever get stuck, there are real good worked out examples here http://www.purplemath.com/modules/hyperbola2.htm

OpenStudy (anonymous):

and if you need to cheat, you can check your answers here http://www.wolframalpha.com/input/?i=hyperbola++%28%28x-5%29^2%2F81%29-%28%28y-1%29^2%29%2F144%29%3D1